# Linking emitted drops to collective bursting bubbles across a wide range of bubble size distributions

### v1

@@ 2.2 Experimental protocol and statistically steady-state measurements of bubbles, drops, and particles
Once the artificial seawater solution is in the tank, a lid is placed on top to seal the system. A constant stream of filtered, compressed air is then flowed through the air above the water surface to flush aerosols from the system. The background aerosols escape through a single port in the lid, which is left open throughout the run to maintain equilibrium within the system. The flushing continues until the CPC component of the SMPS measures a particle count below 10 particles/cm3, as shown in panel (a), which can be compared to an ambient particle count of $\sim$1000 particles/cm3. This ensures that the sea salt aerosols we sample in the air are nearly all emitted by bubble bursting. The air through the top of the tank is then turned off, and the system is left to equilibrate for 40 minutes. CPC particle counts from the pre-run period are shown in panel (b) of figure 3. We confirm that the particle count and relative humidity level (panel (a)) have reached a steady state before proceeding with the run. Relative humidity, air temperature, and water temperature are tracked throughout the run using a Thorlabs sensor (TSP01, TSP-TH) mounted to the tank lid. Values are comparable to realistic ocean conditions and remain steady throughout the run (temperatures vary by $\sim$1°C throughout the measurement period).
@@
Measurement Pixel Size (µm) Field of View (mm2) Recording Framerate (Hz)
bulk 42 80.6x50.4 1
bulk small 5.3 10.6x10.6 2
@@
drops - holography 0.98 3.89x2.19 20
Table 2: Measurement details for imaging of bulk bubbles, surface bubbles, and drops (holography) in the present experiment.
@@ 2.2.1 Details of bulk bubbles
Bulk bubble edge detection was performed using a Canny filter algorithm, which was done in multiple stages for the large field of view to enable successful detection of many sizes of bubbles. An ellipse was fit to each detected region, and the bubble radius $R_{b}$ represents the volume-equivalent size of the bulk bubble, assuming axisymmetry around the minor axis of the ellipse. The void fraction $\phi_{b}$ is typically between 0.1-0.5% for all cases studied, indicating a dilute region where underwater coalescence events are rare.
The depth of field for all cases is calibrated to account for size dependence, using the method detailed in Erinin et al., (2023). For the monodisperse needle cases (where the bubbles rise from four needle arrays and are not evenly distributed throughout the depth), a larger effective depth of field is calculated so that the normalized distribution is appropriately representative of the volume of rising bubbles.
@@ 2.2.2 Details of surface bubbles
At the surface, measurements of the bubbles are again made in two overlapping fields of view for the same range of bubble radii, $R_{b}$ = [30, 5000] µm. The camera for the large field of view (Basler a2A5328-15) captures clusters of bubbles, while the small field of view camera (Basler acA2040-90 with Infinity K2 DistaMax microscope lens) zooms in on the smallest bubbles that rise to the surface. The large and small surface measurements were made at 2 Hz and 3 Hz, respectively, and the frame independence of both measurements was checked by downsampling in time. Details are provided in table 2.
Surface bubbles are detected using the Hough Transform with custom filtering to remove duplicates and misdetection, which enabled detection of both clustered and isolated bubbles within the same frame using a single processing algorithm. The apparent radius of the detected bubbles from the surface measurement is then converted to the volume-equivalent radius $R_{b}$ by considering the static shape of the bubble at the surface, as described in Néel and Deike, (2021), based on the work of Toba, (1959).
@@ 2.2.3 Details of drops and dry particles
As the bubbles burst at the surface, they release drops into the air above the water surface. Three techniques are combined to measure ejected drops with radii spanning four orders of magnitude, in the range $r_{d}$ = [0.05, 500] µm. The first measurement is the digital inline holographic imaging setup (laser: CrystalLaser Nd:YLF QL527-200-L, camera: Phantom VEO4K-990-L, lens: Infinity K2 DistaMax, fully described in Erinin et al., (2023)). The holographic setup is positioned 5.5 cm above the water surface and captures the largest sizes of drops, in the range $r_{d}$ = [10, 500] µm. The small field of view and framerate of 20 Hz are chosen so that each drop is counted only once, either on the upward or downward part of its trajectory.
@@
To quantify the background noise of aerosols remaining in the system, SMPS and OPS measurements were performed for a solution of DI water immediately before each artificial seawater case is run. The calculated background is then subtracted from the size distribution, smoothed using a Savitsky-Golay filter, and cut to the trusted measurement region. In order for all three measurements to be plotted together, the native units output by the SMPS are converted to the normalized concentration $N_{d}(r_{d})$, such that integrating the distribution gives a number of particles per unit volume (see Erinin et al., (2023) for conversion details).
@@ 2.2.4 Typical bubble and drop size distributions
Figure 4: Size distributions of bulk bubbles ($N_{b}(R_{b})$, (a, d)), surface bubbles ($N_{s}(R_{b})$, (b, e)), and drops ($N_{d}(r_{d})$, (c, f)) for broad-banded (a-c) and narrow-banded (d-f) distributions of rising bubbles (see cases Broad-2 mm and Narrow-2 mm and in table 1). Different symbols indicate the different fields of view (bubble measurements) or the different measurement techniques (drop/particle measurements). All distributions are interpolated over the logarithmically-spaced bins, with the interpolated distribution plotted over the data points.
@@
The systematic measurements described above for a wide range of scales allow us to extract full size distributions for both drops and bubbles and probe different conditions in a systematic way (conditions summarized in table 1, with sample surface images in figure 1). We illustrate the characterization of the size distribution for two representative cases in figure 4, the broad and narrow 2 mm cases.
@@
Using bubble sizes and counts extracted from the images through the methods described in the previous sections, we plot the size distributions of the detected bulk and surface bubbles, $N_{b}(R_{b})$ and $N_{s}(R_{b})$, shown in figure 4 normalized by the bin size and by the measurement volume or area, respectively. For both bubble size distributions, data from the large and small field of view measurements are represented by different symbols and nicely overlap. Note that the bubble radius $R_{b}$ represents the volumetric bubble radius for both the bulk and surface measurements.
@@
For the drop size distribution, all measurements are expressed as $N_{d}(r_{d})$, the number of drops per unit bin size and unit measurement volume. Data from the SMPS, OPS, and holographic imaging overlap well, and the combined measurements result in a single distribution that characterizes drops spanning 50 nm to 500 µm.
@@
For the broad-banded case, we observe a continuum of bubble radii spanning $R_{b}$ = [30, 5000] µm in both the bulk and the surface (a, b). We see a corresponding continuum of drop radii spanning $r_{d}$ = [0.05, 500] µm (c). In the narrow-banded case, we see a narrow peak of bubble sizes around the injection bubble radius, $R_{b}$ = 2.3 mm (d, e). We refer to this case as narrow-banded, or nearly monodisperse, because the majority of bubbles are found in this peak around 2 mm. Some amount of smaller bubbles are also present for this narrow-banded case (formed as bubbles pinch off from the needles), and there are also smaller drops, shown in the drop size distribution below $r_{d}\sim$ 200 µm (f). The bulk and surface distributions show similar trends in both conditions, suggesting limited coalescence (which is expected for artificial seawater). Note that some drops above $r_{d}$ = 200 µm may not be detected, either as a result of the small field of view or because they fall back to the surface before reaching the holographic measurement region (Néel and Deike,, 2022). Measurement of drops larger than $r_{d}$ = 600 µm would require a wider field of view, which would limit our ability to measure small drops and connect the holographic measurement to the OPS measurement.
@@ 3 Linking bulk and surface bubble size distributions
The bubble rise velocity $w_{b}(R_{b})$ is a function of the bubble size, and we use the semi-empirical rise velocity parameterization for contaminated water presented in Clift et al., (1978), p. 172. The timescale required for the flux of surface bubbles is the bubble lifetime $\tau_{s}$. Significant uncertainty remains surrounding the dependence of bubble lifetime on the bubble size and properties of the solution (see table 1 in Poulain et al., (2018)), so for simplicity we consider a reasonable estimate of a constant bubble lifetime of order O(1) second (we assume $\tau_{s}=0.6$ s). Equation 1 effectively neglects the effect of coalescence at the surface, which is generally an appropriate simplification given the seawater salinity of the solution (effect of salts on bubble coalescence reviewed in Firouzi et al.,, 2015).
@@
Some words of caution are required when presenting the flux-based comparison. We have chosen one particular rise velocity parameterization and a constant bubble lifetime to evaluate the flux argument. Other relations for the rise velocity were also tested, including the parameterization from Woolf and Thorpe, (1991) which gives a similar result aside from a slightly different velocity for the smallest bubbles. We could also have considered the effect of the background turbulence on the rise velocity, which systematically slows down the rise velocity of the bubbles (Ruth et al.,, 2021) (we note that a systematically lower rise velocity could be compensated for by a slightly higher value of the bubble lifetime, within the uncertainties of experimentally-measured bubble lifetime). We also note that by assuming equation 1 and a rise velocity function, we could also have extracted an effective bubble lifetime function. While uncertainty remains in both the chosen rise velocity and lifetime, we find that we can relate the bulk and surface bubble size distributions well for all the cases presented in the paper, so that we can use the bulk and surface measurements interchangeably by applying equation 1.
In the following analysis, we will use an effective size distribution for bursting bubbles (using both surface and bulk data) to practically relate the bubbles to the drop size distribution, which is measured per unit volume. A direct link between bursting bubbles and emitted drops in terms of fluxes would require measurements of the drop velocity (not available here). To come up with this effective bubble size distribution, the surface bubble distribution is first converted to a quantity of bursting bubbles normalized by a volume, following equation 1. For bubbles below $R_{b}$ = 150 µm, we use data from the bulk small field of view measurement to construct a complete size distribution of bubbles spanning $R_{b}$ = [30, 5000] µm. Throughout the rest of the paper, the term $N_{b}(R_{b})$ is used to represent this size distribution of bursting bubbles from the combined measurements, with the exception of the Narrow-250 mm case, which uses only the converted surface data due to imaging/detection difficulties for bubbles in the denser bulk plume.
@@ 4.1 Direct attribution between measured bubbles and drops
Figure 6: Size distributions of bursting bubbles and drops for various case comparisons. The bubble size distributions presented vary significantly, including broad-banded cases and three narrow-banded cases peaking at radii of 2 mm, 3 mm, and 250 µm, with a corresponding variety of drop size distributions. To calculate the drop production efficiency presented in the rightmost column, the bubble and drop size distributions are first integrated in the shaded or outlined size buckets, where the link between the bubble and drop sizes approximates the link suggested by individual bubble scaling laws. The resulting number of drops in each size range is then divided by the number of bubbles in the corresponding shaded size range (outlined for film flap range) to obtain an average number of drops produced per bubble in each bubble size bucket. Shaded grey lines show the number of drops per bubble predicted by the individual bubble bursting scalings for jet (solid line) and film flapping (dotted line) production mechanisms (tested in section 4.2, see table 3 for details). Cases shown are (a-c): Narrow-2 mm vs. Broad-2 mm, (d-f): Narrow-3 mm vs. Narrow-250 µm, and (g-i): Broad-2 mm vs. Broad-half volume. Further details of each case are found in table 1.
@@
The first row (a-c) of figure 6 shows the two sample cases described in section 2.2.4 (figures 4 and 5, narrow and broad 2 mm injection cases). Both cases use the same needles and flow rate, but in the broad case turbulent flow is created underwater. As described above, turbulence induces bubble breakup and creates a broad-banded bubble size distribution with a continuum of bubbles spanning radii of 30 µm to almost 5 mm. Without turbulence, the distribution has a peak around $R_{b}$ = 2 mm with a lower number of smaller bubbles. The two cases (broad- and narrow-banded) have a relatively comparable number of bubbles around 2 mm, near the needle injection size. For the submillimetric bubbles, there is a large difference of about two orders of magnitude in number of bubbles (because without turbulence, the only small bubbles present are formed by pinch-off at the needles, see section 2). Moving to the drop size distribution, we observe a corresponding difference of similar magnitude between the drop size distributions for the entire range of sizes measured in the experiment. This suggests that the drops we measure below radius 200 µm are produced by bubbles with radii smaller than $\sim$2 mm (shaded in color).
@@
Size buckets are chosen to approximate the individual bubble size scaling laws for jet and film flapping production mechanisms (discussed in more detail in section 4.2, and Berny et al., (2021); Jiang et al., (2022); Deike et al., (2022)). The corresponding bubble and drop ranges are shown in shaded/outlined colors in figure 6, where the overlap in bubble sizes represents the fact that bubbles can produce both jet and film flap drops in parallel processes. For the film flap scaling, we test the link between bubbles of radii $R_{b}$ = [70,1200] µm and drops of radii $r_{d}$ $<$ 0.6 µm. For the jet drop scaling, we consider the ability of bubbles in the range $R_{b}$ = [30, 1500] µm to produce drops of $r_{d}$ $>$ 0.6 µm. The total regions tested for jet drop production are evenly divided into 3 smaller log-spaced size buckets for both bubbles and drops, and the buckets are linked in size order.
@@
Resulting drop production efficiencies are shown in the right column of figure 6 (c, f, i), calculated from the shaded ranges (jet drop) or outlined ranges (film flap drop) of bubble and drop size buckets for the two cases. Note that the values of the efficiencies would shift somewhat if we chose to split the drop size distribution at a value other than $r_{d}$ = 0.6 µm when testing the attribution to the assumed bubble size ranges. The cutoff radius at $r_{d}$ = 0.6 µm was chosen because it falls between the largest predicted film drop and smallest predicted jet drop for the range of bubble sizes present in the experiment.
For the broad-banded case, we observe $\sim$50 drops/bubble in the assumed film flap range (c, black diamond) and $\sim$0.8–1.5 drops/bubble across the jet drop size buckets (c, black circles). For the narrow-banded case, we observe a high efficiency of around 100 drops/bubble in the film flap size range (c, blue diamond), while the small jet drop efficiency is around 10 and the larger jet drop efficiency is below unity (c, blue circles).
These numbers can be compared to the efficiencies discussed in the literature. The light gray lines in the same plot represent the drop number given by the individual bubble bursting scaling laws used throughout the paper for jet drops (solid, between 0.6–3 drops produced per bubble, Wang et al., (2017)) and film flap drops (dotted, 40 drops per bubble, Jiang et al., (2022)). Detailed equations for the scalings shown here, based on past work in the literature, are provided in table 3 and described in section 4.2.
Comparing the calculated efficiencies for the two sample cases to the approximate individual bursting scaling laws, we see reasonable agreement for the broad-banded case. For the nearly monodisperse case, significant differences are observed, and various interpretations can be proposed. The higher efficiency of film flap drops could be related to the proposed bubble pinch-off effect/small aerosol formation in the bulk discussed in the recent work from Jiang et al., (2024), with this effect potentially smeared out in the presence of turbulence. The higher jet drop efficiency for the small bubbles, which closely match direct numerical simulations for similar conditions (Berny et al.,, 2021), could be due to the fact that either low surface density of bubbles or a calmer free surface are necessary to obtain such efficiency. Larger jet drop low production efficiency might be related to a lack of statistical convergence in measuring large drops (related to the small field of view and height effect).
@@
From the size distributions, it is immediately evident that many more submicron droplets are produced (e) when more small submillimetric bubbles are present (d). We observe an approximately 4 order-of-magnitude difference between the size distributions of the small bubble case and the large bubble case, where very few submicron drops are produced. As shown by the large bubble case, when we measure very few submillimetric bubbles, we measure very few submicron drops. This demonstrates that bubbles predominantly larger than 2 mm do not produce a significant amount of submicron aerosols in this configuration.
@@
Comparing to individual bursting scalings, we see a reasonable efficiency of $\sim$60 drops/bubble for both cases for the proposed film flap range (f, green and orange diamonds). In the two largest jet drop bins, the efficiencies for bubbles are low (f, green and orange circles; efficiencies below 0.1 associated with very small number of drops are not shown on the plot) with similar values to those discussed above for the 2 mm narrow-banded case (c, blue circles). For the smallest jet drop bin, the efficiencies are high, at around 20 drops/bubble. We note that the overall efficiencies are consistent across all the cases despite the large differences in the bubble size distributions (see the plot of number of bubbles versus number of drops in appendix A), and it is only once we assume a link between the bubbles sizes and drops sizes for this attribution that differences in efficiencies appear for the narrow-banded cases compared to the broad-banded cases. The different efficiency values found in the narrow-banded cases imply that some subtleties of the bubble-drop link may be blurred across neighboring size buckets for the continuum of bubble and drops sizes present in the broad-banded cases, which seem to match the single bubble scaling laws well.

### v2

@@ 2.2 Experimental protocol and statistically steady-state measurements of bubbles, drops, and particles
Once the artificial seawater solution is in the tank, a lid is placed on top to seal the system. A constant stream of filtered, compressed air is then flowed through the air above the water surface to flush aerosols from the system. The background aerosols escape through a single port in the lid, which is left open throughout the run to maintain equilibrium within the system. The flushing continues until the CPC component of the SMPS measures a particle count below 10 particles/cm3, as shown in panel (b), which can be compared to an ambient particle count of $\sim$1000 particles/cm3. This ensures that the sea salt aerosols we sample in the air are nearly all emitted by bubble bursting. The air through the top of the tank is then turned off, and the system is left to equilibrate for 40 minutes. CPC particle counts from the pre-run period are shown in panel (b) of figure 3. We confirm that the particle count and relative humidity level (panel (a)) have reached a steady state before proceeding with the run. Relative humidity, air temperature, and water temperature are tracked throughout the run using a Thorlabs sensor (TSP01, TSP-TH) mounted to the tank lid. Values are comparable to realistic ocean conditions and remain steady throughout the run (temperatures vary by $\sim$1°C throughout the measurement period).
@@
Measurement Pixel Size (µm) Field of View (mm2) Depth of Field (mm) Recording Framerate (Hz)
bulk 42 80.6 x 50.4 14–26 1
bulk small 5.3 10.6 x 10.6 2 2
@@
drops - holography 0.98 3.89 x 2.19 360 20
drops - shadow 29 59.4 x 59.4 120 5
Table 2: Measurement details for imaging of bulk bubbles, surface bubbles, and drops (holography and shadow) in the present experiment. The shadow imaging drop measurement was performed for certain cases (Narrow - 20mm, Narrow - 30mm, Narrow - 2500µm, and Broad - 20mm) to allow for the measurement of drops very close to the surface. A range is specified for the depth of field of the bulk measurement resulting from the size-dependent depth-of-field correction, and this range corresponds to the depth of field of bubble radii in the range $R_{b}$ = [150, 3000]0µm (see details in section 2.2.3 and figure 6 of Erinin et al. (2023)).
@@ 2.2.1 Details of bulk bubbles
Bulk bubble edge detection was performed using a Canny filter algorithm, which was done in multiple stages for the large field of view to enable successful detection of many sizes of bubbles. An ellipse was fit to each detected region, and the bubble radius $R_{b}$ represents the volume-equivalent size of the bulk bubble, assuming axisymmetry around the minor axis of the ellipse. The void fraction $\phi_{b}$ is typically between 0.1-0.5% for all cases studied, indicating a dilute region of bubbly flow. Through inspection of dedicated high-speed movies, we confirm that underwater collisions and coalescence in the bulk are very rare. The presence of salt limits the coalescence efficiency so that if bubbles do contact each other, they typically bounce off and do not coalesce.
The depth of field for all cases is calibrated to account for its dependence on the size of the bubble, using the method detailed in section 2.2.3 and figure 6 of Erinin et al. (2023). For the Narrow - 2 mm and Narrow - 3 mm cases, the depth of field is additionally adjusted to account for the fact that the bulk bubbles are not evenly distributed throughout the depth (because they are rising from a square array of needles). This adjustment leads to a larger effective depth of field for the narrow-banded cases, which enables comparison to the broad-banded cases where the bubbles are distributed evenly throughout the volume by the underwater pumps. The average relative error on the bubble radius measurement is 6% at the limits of the depth of field, see Erinin et al. (2023) (section 2.2.3 and figure 7).
@@ 2.2.2 Details of surface bubbles
At the surface, measurements of the bubbles are again made in two overlapping fields of view for the same range of bubble radii, $R_{b}$ = [30, 5000] µm. The camera for the large field of view (Basler a2A5328-15) captures clusters of bubbles, while the small field-of-view camera (Basler acA2040-90 with Infinity K2 DistaMax microscope lens) zooms in on the smallest bubbles that rise to the surface. The large and small surface measurements were made at 2 Hz and 3 Hz, respectively, and the frame independence of both measurements was checked by downsampling in time. Details are provided in table 2.
Surface bubbles are detected using the Hough Transform with custom filtering to remove duplicates and misdetection, which enabled detection of both clustered and isolated bubbles within the same frame using a single processing algorithm. The apparent radius of the detected bubbles from the surface measurement is then converted to the volume-equivalent radius $R_{b}$ by considering the static shape of the bubble at the surface, as described in Néel & Deike (2021), based on the work of Toba (1959).
@@ 2.2.3 Details of drops and dry particles
As the bubbles burst at the surface, they release drops into the air above the water surface. Three techniques are combined to measure ejected drops with radii spanning four orders of magnitude, in the range $r_{d}$ = [0.05, 500] µm. The first measurement is the digital inline holographic imaging setup (laser: CrystalLaser Nd:YLF QL527-200-L, camera: Phantom VEO4K-990-L, lens: Infinity K2 DistaMax, fully described in Erinin et al. (2023)). The holographic setup is positioned 5.5 cm above the water surface and captures the large sizes of drops, in the range $r_{d}$ = [10, 500] µm. The small field of view and framerate of 20 Hz are chosen so that each drop is counted only once, either on the upward or downward part of its trajectory.
@@
To quantify the background noise of aerosols remaining in the system, SMPS and OPS measurements were performed for a solution of DI water immediately before each artificial seawater case is run. The calculated background is then subtracted from the size distribution, smoothed using a Savitsky-Golay filter, and cut to the trusted measurement region. In order for all three measurements to be plotted together, the native units output by the SMPS are converted to the normalized concentration $N_{d}(r_{d})$, such that integrating the distribution gives a number of particles per unit volume (see Erinin et al. (2023) for conversion details).
We note that some drops above $r_{d}$ = 200 µm may not be detected by the holographic measurement because they fall back to the surface before reaching the measurement region (Néel & Deike, 2022). To confirm the presence of these large drops, the drop measurements were supplemented with a shadow imaging measurement (for drops of radii $r_{d}>$ 150 µm) for the Narrow-2 mm, Narrow-3 mm, Narrow-250 µm, and Broad-2 mm cases. This measurement was added directly above the level of the water surface to quantify large drops that may not have reached the holographic measurement region (camera: Basler acA2040-90, telecentric lens: Opto-E TC4MHR096-C). Details of the shadow measurement are included in table 2, and the imaging setup is described for drop measurements in Erinin et al. (2023).
@@ 2.2.4 Typical bubble and drop size distributions
Figure 4: Size distributions of bulk bubbles ($N_{b}(R_{b})$, (a, d)), surface bubbles ($N_{s}(R_{b})$, (b, e)), and drops ($N_{d}(r_{d})$, (c, f)) for broad-banded (a-c) and narrow-banded (d-f) distributions of rising bubbles (see cases Broad-2 mm and Narrow-2 mm and in table 1). Different symbols indicate the different fields of view (bubble measurements) or the different measurement techniques (drop/particle measurements). The drop size distribution is shown in terms of the liquid drop radius, $r_{d}$. All distributions are interpolated over the logarithmically-spaced bins, with the interpolated distribution plotted over the data points.
@@
The systematic measurements described above for a wide range of scales allow us to extract full size distributions for both drops and bubbles and probe different conditions in a systematic way (conditions summarized in table 1, with sample surface images in figure 1). We illustrate the characterization of the size distribution for two representative cases in figure 4, the Broad-2 mm and Narrow-2 mm cases.
@@
Using bubble sizes and counts extracted from the images through the methods described in the previous sections, we plot the size distributions of the detected bulk and surface bubbles, $N_{b}(R_{b})$ and $N_{s}(R_{b})$, shown in figure 4 normalized by the bin size and by the measurement volume or area, respectively. For both bubble size distributions, data from the large and small field-of-view measurements are represented by different symbols and nicely overlap. Note that the bubble radius $R_{b}$ represents the volumetric bubble radius for both the bulk and surface measurements.
@@
For the drop size distribution, all measurements are expressed as $N_{d}(r_{d})$, the number of drops per unit bin size and unit measurement volume. Data from the SMPS, OPS, holographic imaging, and shadow imaging overlap well, and the combined measurements result in a single distribution that characterizes drops spanning 50 nm to 500 µm.
@@
For the broad-banded case, we observe a continuum of bubble radii spanning $R_{b}$ = [30, 5000] µm in both the bulk and the surface (a, b). We see a corresponding continuum of drop radii spanning $r_{d}$ = [0.05, 500] µm (c). In the narrow-banded case, we see a narrow peak of bubble sizes around the injection bubble radius, $R_{b}$ = 2.3 mm (d, e). We refer to this case as narrow-banded, or nearly monodisperse, because the majority of bubbles are found in this peak around 2 mm. Some amount of smaller bubbles are also present for this narrow-banded case (formed as bubbles pinch off from the needles), and there are also smaller drops, shown in the drop size distribution below $r_{d}\sim$ 200 µm (f). The bulk and surface distributions show similar trends in both conditions, suggesting limited coalescence (which is expected for artificial seawater).
@@ 3 Linking bulk and surface bubble size distributions
The bubble rise velocity $w_{b}(R_{b})$ is a function of the bubble size, and we use the semi-empirical rise velocity parameterization for contaminated water presented in Clift et al. (1978), p. 172. The timescale required for the flux of surface bubbles is the bubble lifetime $\tau_{s}$. Significant uncertainty remains surrounding the dependence of bubble lifetime on the bubble size and properties of the solution (see table 1 in Poulain et al. (2018)), so for simplicity we consider a reasonable estimate of a constant bubble lifetime of order O(1) second (we assume $\tau_{s}=0.6$ s). Equation 1 effectively neglects the effect of coalescence at the surface, which is generally an appropriate simplification given the seawater salinity of the solution (effect of salts on bubble coalescence reviewed in Firouzi et al., 2015).
@@
[…]
In the following analysis, we will use an effective size distribution for bursting bubbles (using both surface and bulk data) to practically relate the bubbles to the drop size distribution, which is measured per unit volume. A direct link between bursting bubbles and emitted drops in terms of fluxes would require measurements of the drop velocity (not available here). To come up with this effective bubble size distribution, the surface bubble distribution is first converted to a quantity of bursting bubbles normalized by a volume, following equation 1. For bubbles below $R_{b}$ = 150 µm, we use data from the bulk small field-of-view measurement to construct a complete size distribution of bubbles spanning $R_{b}$ = [30, 5000] µm. Throughout the rest of the paper, the term $N_{b}(R_{b})$ is used to represent this size distribution of bursting bubbles from the combined measurements, with the exception of the Narrow-250 µm case, which uses only the converted surface data due to imaging/detection difficulties for bubbles in the denser bulk plume.
@@ 4.1 Direct attribution between measured bubbles and drops
Figure 6: Size distributions of bursting bubbles ($N_{b}(R_{b})$, (a, d, g)) and drops ($N_{d}(r_{d})$, (b, e, h), in terms of the liquid drop radius, $r_{d}$) for various case comparisons. The bubble size distributions presented vary significantly, including broad-banded cases and three narrow-banded cases peaking at radii of 2 mm, 3 mm, and 250 µm, with a corresponding variety of drop size distributions. To calculate the drop production efficiency presented in the rightmost column (c, f, i), the bubble and drop size distributions are first integrated in the shaded or outlined size buckets, where the link between the bubble and drop sizes approximates the link suggested by scaling laws for individual bubble bursting. The resulting number of drops in each size range is then divided by the number of bubbles in the corresponding shaded size range (outlined for film flap range) to obtain an average number of drops produced per bubble in each bubble size bucket (production efficiency $\eta_{d}$, shown as circles within the jet drop size range corresponding to the blue shaded regions and diamonds for the assumed film flap drop size range corresponding to the orange boxes). Shaded gray lines show the number of drops per bubble predicted by the individual bubble bursting scalings for jet (solid line) and film flapping (dotted line) production mechanisms (tested in section 4.2, see table 3 for details). Cases shown are (a-c): Narrow-2 mm vs. Broad-2 mm, (d-f): Narrow-3 mm vs. Narrow-250 µm, and (g-i): Broad-2 mm vs. Broad-2 mm-half volume. Further details of each case are found in table 1, and the uncertainty associated with the production efficiency values shown in (c, f, i) is discussed in appendix B.
@@
The first row (a-c) of figure 6 shows the two sample cases described in section 2.2.4 (figures 4 and 5, Narrow and Broad 2 mm injection cases). Both cases use the same needles and flow rate, but in the broad case turbulent flow is created underwater. As described above, turbulence induces bubble breakup and creates a broad-banded bubble size distribution with a continuum of bubbles spanning radii of 30 µm to almost 5 mm. Without turbulence, the distribution has a peak around $R_{b}$ = 2 mm with a lower number of smaller bubbles. The two cases (broad- and narrow-banded) have a relatively comparable number of bubbles around 2 mm, near the needle injection size. For the submillimetric bubbles, there is a large difference of about two orders of magnitude in number of bubbles (because without turbulence, the only small bubbles present are formed by pinch-off at the needles, see section 2). Moving to the drop size distribution, we observe a corresponding difference of similar magnitude between the drop size distributions for nearly the entire range of sizes measured in the experiment, until the number of drops becomes more comparable between cases for $r_{d}>$ 400 µm. This suggests that the drops we measure below radius 400 µm are produced by bubbles with radii smaller than $\sim$2 mm (shaded in color).
@@
Size buckets are chosen to approximate the individual bubble size scaling laws for jet and film flapping production mechanisms (discussed in more detail in section 4.2, and Berny et al. (2021); Jiang et al. (2022); Deike et al. (2022)). The corresponding bubble and drop ranges are shown in shaded/outlined colors in figure 6, where the overlap in bubble sizes represents the fact that bubbles can produce both jet and film flap drops in parallel processes. For the film flap scaling, we test the link between bubbles of radii $R_{b}$ = [70, 1200] µm and drops of radii $r_{d}$ $<$ 0.5 µm (outlined/shaded in orange). For the jet drop scaling, we consider the ability of bubbles in the range $R_{b}$ = [30, 2000] µm to produce drops of $r_{d}$ $>$ 0.5 µm (shaded in blue). The total regions tested for jet drop production are approximately evenly divided into 3 smaller log-spaced size buckets for both bubbles and drops, and the buckets are linked in size order.
@@
Resulting drop production efficiencies are shown in the right column of figure 6 (c, f, i), calculated from the shaded ranges (jet drop, circles) or outlined ranges (film flap drop, diamonds) of bubble and drop size buckets for the two cases. Note that the values of the efficiencies would shift somewhat if we chose to split the drop size distribution at a value other than $r_{d}$ = 0.5 µm when testing the attribution to the assumed bubble size ranges. The cutoff radius at $r_{d}$ = 0.5 µm was chosen because it falls between the largest predicted film drop and smallest predicted jet drop for the range of bubble sizes present in the experiment.
For the broad-banded case, we observe $\sim$50 drops/bubble in the assumed film flap range (c, black diamond) and $\sim$1.0–1.8 drops/bubble across the jet drop size buckets (c, black circles). For the narrow-banded case, we observe a high efficiency of around 100 drops/bubble in the film flap size range (c, blue diamond), while the small jet drop efficiency is around 9 and the larger jet drop efficiency is below unity (c, blue circles).
These numbers can be compared to the efficiencies discussed in the literature. The light gray lines in the same plot represent the drop number given by the individual bubble bursting scaling laws used throughout the paper for jet drops (solid, between 0.6–3 drops produced per bubble, Wang et al. (2017)) and film flap drops (dotted, 40 drops per bubble, Jiang et al. (2022)). Detailed equations for the scalings shown here, based on past work in the literature, are provided in table 3 and described in section 4.2.
Comparing the calculated efficiencies for the two sample cases to the approximate individual bursting scaling laws, we see reasonable agreement for the broad-banded case. For the nearly monodisperse case, significant differences are observed, and various interpretations can be proposed. The higher efficiency of film flap drops could be related to the proposed bubble pinch-off effect/small aerosol formation in the bulk discussed in the recent work from Jiang et al. (2024), with this effect potentially smeared out or not occurring in the presence of turbulence (note that for the Narrow-2 mm case, high-speed movies confirm collisions at the needle tip/during bubble pinch-off as discussed in Jiang et al. (2024). These collisions at the needle tips are not observed in the Narrow-3 mm case). The higher jet drop efficiency for the small bubbles, which closely match direct numerical simulations for similar conditions (Berny et al., 2021), could be due to the fact that either low surface density of bubbles or a calmer free surface are necessary to obtain such efficiency. Larger jet drop low production efficiency might be related to a lack of statistical convergence in measuring these drops with the holographic system. Sources of uncertainty involved in the calculation of the production efficiency values presented in figure 6 are discussed in appendix B.
@@
From the size distributions, it is immediately evident that many more submicron droplets are produced (e) when more small submillimetric bubbles are present (d). We observe an approximately 3 order-of-magnitude difference between the size distributions of the small bubble case and the large bubble case, where very few submicron drops are produced. As shown by the large bubble case, when we measure very few submillimetric bubbles, we measure very few submicron drops. This demonstrates that bubbles predominantly larger than 2 mm do not produce a significant amount of submicron aerosols in this configuration.
@@
Comparing to individual bursting scalings, we see a reasonable efficiency of $\sim$60 drops/bubble for both cases for the proposed film flap range (f, green and orange diamonds). In the two largest jet drop bins, the efficiencies for bubbles are low (f, green and orange circles; efficiencies below 0.1 associated with very small number of drops are not shown on the plot) with similar values to those discussed above for the 2 mm narrow-banded case (c, blue circles). For the smallest jet drop bin, the efficiencies are high, at around 20 drops/bubble. We note that the overall efficiencies are consistent across all the cases despite the large differences in the bubble size distributions (see the plot of number of drops versus number of bubbles in appendix A), and it is only once we assume a link between the bubbles sizes and drop sizes for this attribution that differences in efficiencies appear for the narrow-banded cases compared to the broad-banded cases. The different efficiency values found in the narrow-banded cases imply that some subtleties of the bubble-drop link may be blurred across neighboring size buckets for the continuum of bubble and drop sizes present in the broad-banded cases, which seem to match the single bubble scaling laws well.